# Neuronal correlations shape the scaling behavior of memory capacity and nonlinear computational capability of recurrent neural networks

### v1

@@ Acknowledgements.
We thank Joni Dambre for providing the source code for computing information processing capacity. S.T. was supported by JSPS KAKENHI Grant No. JP22KJ1959. T.A. was supported by MEXT KAKENHI Grant Numbers 23H04467 and by JSPS KAKENHI Grant Numbers 24H00723, 20K20520.
@@ B.1 Analytical derivation of memory capacity
[…]
$MC=\sum_{d=0}^{\infty}\sum_{n=0}^{\infty}(-1)^{n}\left\{\frac{\alpha\tilde{\sigma}_{s}^{2}}{[\langle x_{i}^{2}\rangle]}\left(\frac{g^{2}}{1+\frac{\pi}{2}[\langle x_{i}^{2}\rangle]}\right)^{d}\right\}^{n+1},$ (63)
where $[\langle x_{i}^{2}\rangle]$ is determined by solving
$[\langle x_{i}^{2}\rangle]=\sigma_{n}^{2}+g^{2}\left(-1+\frac{4}{\pi}\arctan\sqrt{1+\pi[\langle x_{i}^{2}\rangle]}\right).$ (64)
@@ B.2 Analytical solution of the decay rate, $\lim_{N\to\infty}r(L)$
Fig.6 illustrates the dependence of the theoretically derived decay rate on the model parameters, $\tilde{\sigma}_{s}$, $g$, and $\sigma_{n}$. As can be seen, the decay rate decreases rapidly for large inputs ($\tilde{\sigma}_{s}$), small recurrent weights ($g$), and small neuronal noise ($\sigma_{n}$). Notably, these changes in hyperparameters result in increased neuronal correlations, which can be confirmed by Eq.(38). Consequently, we can expect the second term in Eq.(69), which originates from neuronal correlations, to grow as the neuronal correlations increase, leading to a faster decline in the decay rate.
[…]
@@ C.1 Overview of IPC theory
We provide a concise overview of the IPC, including its definition and key properties. We consider a reservoir receiving input signals $\{s(t)\}_{t}$. The general task is formulated as optimizing output weights such that the reservoir’s output $\hat{z}(t)$ approximates a given function of the inputs, $f[\cdots,s(t-1),s(t)]$. For input signals drawn i.i.d from a standard Gaussian distribution, the orthonormal basis functions spanning the Hilbert space containing $f$ are expressed as a formally infinite product of normalized Hermite polynomials:
@@
$C_{T}[f]\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(\hat{z}(t)-f(t))^{2}\rangle_{T}}{\langle f(t)^{2}\rangle_{T}},$ (72)

### v2

@@ Acknowledgements.
We thank Joni Dambre for providing the source code for computing information processing capacity, and Tomoyuki Kubota for helpful discussions regarding its implementation. S.T. was supported by JSPS KAKENHI Grant No. JP22KJ1959. T.A. was supported by MEXT KAKENHI Grant Numbers 23H04467 and by JSPS KAKENHI Grant Numbers 24H00723, 20K20520.
@@ C.1 Overview of IPC theory
We provide a concise overview of the IPC, including its definition and key properties. We consider a reservoir receiving input signals $\{s(t)\}_{t}$. The general task is formulated as optimizing output weights such that the reservoir’s output $z(t)$ approximates a given function of the inputs, $f[\cdots,s(t-1),s(t)]$. For input signals drawn i.i.d from a standard Gaussian distribution, the orthonormal basis functions spanning the Hilbert space containing $f$ are expressed as a formally infinite product of normalized Hermite polynomials:
@@
$C_{T}[f]\equiv 1-\frac{{\rm{min}}_{\bm{w}}\langle(z(t)-f(t))^{2}\rangle_{T}}{\langle f(t)^{2}\rangle_{T}},$ (73)